# Congestion-Approximators from the Bottom Up

### v2

@@ A.2.2 Max-flow/min-cut call
The path-removing step in the algorithm is then possible, and ensures that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of exactly $\mathbf{d}(v)/12$ flow in $f_{t}$, each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow, and there are no (nonzero) sources or sinks elsewhere.
We begin with the unscaled flow $f$. Since $(S,f)$ is a $(1+\epsilon)$-fair cut/flow pair, each vertex $v\in V_{H}\setminus(S\cup\{t\})$ receives at least $\frac{1}{1+\epsilon}\,c_{H}(\{v\},S)\geq\frac{1}{2}c_{H}(\{v\},S)$ total flow from vertices in $S$. Since $A\setminus S_{t}\subseteq V_{H}\setminus(S\cup\{t\})$, the same holds for all $v\in A\setminus S_{t}$. By construction of $H$, we have $c_{H}(\{v\},S)\geq c_{H}(v,s)=\phi\mathbf{d}|_{A_{t}^{\ell}}(v)+\epsilon\phi\mathbf{d}(v)$ for all $v\in A\setminus S_{t}$. It follows that each vertex $v\in A\setminus S_{t}$ receives at least $\frac{\phi}{2}\mathbf{d}|_{A_{t}^{\ell}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ total flow from vertices in $S$.
We now investigate the effect of restricting the flow $f$ to edges in $G[A\setminus S_{t}]$, starting with removing all edges incident to $S$. Continuing the argument above, removing these edges causes each vertex $v\in A\setminus S_{t}$ to be the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t}^{\ell}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow.
If $x\notin S$, then we now remove the edges incident to $x$. By construction of $H$, each vertex $v\in A$ has an edge to $x$ of capacity $\gamma\mathbf{d}(v)=\frac{\epsilon\phi}{2}\mathbf{d}(v)$. Since the flow $f$ is feasible, there is at most $\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow along the edge $(v,x)$. Removing this edge changes the net flow out of $v$ by at most $\frac{\epsilon\phi}{2}\mathbf{d}(v)$. Since each vertex $v\in A\setminus S_{t}$ is the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t}^{\ell}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow before this step, it is the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t}^{\ell}}(v)\geq 0$ flow after this step. In particular, it cannot become a sink. Also, if $v\in A_{t}^{\ell}\setminus S_{t}$, then it is the source of at least $\frac{\phi}{2}\mathbf{d}(v)$ flow after restriction.
@@
Finally, scaling this restricted flow by $1/(12\phi)$, we conclude that the scaled flow $f$ has congestion $1/(12\phi)$, each vertex $v\in A_{t}^{\ell}\cap S_{t}$ is the source of at least $\mathbf{d}(v)/12$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. ∎
@@ A.3 Analyzing the Potential Decrease
$\deg_{M_{t}}(u)=\mathbf{d}(u)/48$ for all $u\in A_{t}^{\ell}\setminus S_{t}$, and $\deg_{M_{t}}(v)\leq\mathbf{d}(v)$ for all $u\in A_{t}^{r}\setminus S_{t}$.
@@
For each vertex $u\in A_{t}^{\ell}\setminus S_{t}$, its degree in $M_{t}$ equals the total capacity of paths starting at $u$ in the decomposition of $f_{t}$, which is exactly $\mathbf{d}(u)/48$ by Claim A.8. A symmetric argument establishes $\deg_{M_{t}}(v)\leq\mathbf{d}(v)$ for all $u\in A_{t}^{r}\setminus S_{t}$. ∎
@@
The following lemma adapts Lemma 3.4 of [33] to the capacitated case.
@@
$\psi(t)-\psi(t-1)\geq\displaystyle\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}$.
@@
Subclaim 1.
$0\preceq D^{-1/2}LD^{-1/2}\preceq 2I$.
@@
Consider the normalized Laplacian $D_{L}^{-1/2}LD^{-1/2}_{L}$ where $D_{L}$ is the diagonal matrix with value $\deg_{M_{t}}(u)$ on each entry $(u,u)$. It is well-known that the normalized Laplacian has eigenvalues in the range $[0,2]$, so $0\preceq D_{L}^{-1/2}LD_{L}^{-1/2}\preceq 2I$. Multiplying by $D_{L}^{1/2}$ on both sides gives $0\preceq L\preceq 2D_{L}$. Claim A.14 implies that $D_{L}\preceq D$, so we obtain $0\preceq L\preceq 2D_{L}\preceq 2D$. Finally, multiplying by $D^{-1/2}$ on both sides gives the desired $0\preceq D^{-1/2}LD^{-1/2}\preceq 2I$. ∎
@@
For two matrices $A,B\in\mathbb{R}^{A\times A}$, define the Hadamard product $A\bullet B=\sum_{u,v\in A}A(u,v)\cdot B(u,v)=\textup{Tr}(A^{\top}B)$. To bound $\psi(t-1)$, we use the fact that $\boldsymbol{\mu}_{t-1}$ minimizes the expression in (2) to get
@@
$\leq\sum_{u\in A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t}\right\rVert_{2}^{2}$
$=\sum_{u\in A_{t-1}}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)^{1/2}}-\mathbf{d}(u)^{1/2}\boldsymbol{\mu}_{t}\right\rVert_{2}^{2}$
$=(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}).$
For $\psi(t)$, we use $A_{t-1}\subseteq A_{t}$ to get
@@
$\geq\sum_{u\in A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t}\right\rVert_{2}^{2}$
$=\sum_{u\in A_{t-1}}\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)^{1/2}}-\mathbf{d}(u)^{1/2}\boldsymbol{\mu}_{t}\right\rVert_{2}^{2}$
$=(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}).$
Taking the difference and expanding,
$\psi(t-1)-\psi(t)$
$\geq(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})-(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})$
$=(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})-(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t})$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-2\widetilde{\mathbf{F}}_{t-1}\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}+D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}+2\widetilde{\mathbf{F}}_{t}\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}-D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}+2(\widetilde{\mathbf{F}}_{t}-\widetilde{\mathbf{F}}_{t-1})\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}-2\bigg{(}\frac{1}{2}D^{-1/2}LD^{-1/2}\bigg{)}\bullet D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}.$
The third term equals $\textup{Tr}\big{(}D^{-1/2}LD^{-1/2}D^{1/2}\mathbbm{1}\boldsymbol{\mu}_{t}\big{)}$, which equals 00 since $LD^{-1/2}D^{1/2}\mathbbm{1}=L\mathbbm{1}$ is the zero matrix. Expanding the first and second terms, we obtain
$\psi(t-1)-\psi(t)$ $\geq\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}$
@@
To bound the first term in (5), recall from Subclaim 1 that $0\preceq D^{-1/2}LD^{-1/2}\preceq 2I$, which means that $-I\preceq N\preceq I$. Since $\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top}\succeq 0$ and $N$ has eigenvalues in the range $[-1,1]$, we have $\textup{Tr}(N\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top}N)\leq\textup{Tr}(\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top})$, so
@@
$\psi(t-1)-\psi(t)\geq\frac{1}{2}D^{-1/2}LD^{-1/2}\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}.$
@@
$\psi(t-1)-\psi(t)\geq\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\cdot\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}.\qed$
@@
$\psi(t-1)-\psi(t)\geq\displaystyle\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}.$ (6)
@@
$\mathbb{E}[(\mathbf{p}(u)-\mathbf{p}(v))^{2}]\leq\frac{C\log n}{n}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}$ (8)
for all pairs $u,v\in A$. By Claim A.14, we have $\deg_{M_{t}}(u)=\mathbf{d}(u)/12$ for all $u\in A_{t}^{\ell}\setminus S_{t}$, and together with statements a to c of Lemma A.5, we have
@@
$=\sum_{u\in A_{t}^{\ell}}\deg_{M_{t}}(u)\cdot(\mathbf{p}(u)-\eta)^{2}$
$=\frac{1}{12}\sum_{u\in A_{t}^{\ell}}\mathbf{d}(u)\cdot(\mathbf{p}(u)-\eta)^{2}$
@@
$=\mathbb{E}\bigg{[}\displaystyle\sum_{u\in A_{t}^{\ell},v\in A_{t}^{r}}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}\bigg{]}$
@@
$=\frac{n}{24C\log n}\sum_{u\in A_{t-1}}\mathbf{d}(u)\cdot\mathbb{E}[(\mathbf{p}(u)-\bar{\mu}_{t-1})^{2}]$
@@
$=\frac{1}{24C\log n}\cdot\psi(t-1),$

### v3

@@ A.2.2 Max-flow/min-cut call
The path-removing step in the algorithm is then possible, and ensures that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of exactly $\mathbf{d}(v)/24$ flow in $f_{t}$, each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow, and there are no (nonzero) sources or sinks elsewhere.
We begin with the unscaled flow $f$. Since $(S,f)$ is a $(1+\epsilon)$-fair cut/flow pair, each vertex $v\in V_{H}\setminus(S\cup\{t\})$ receives at least $\frac{1}{1+\epsilon}\,c_{H}(\{v\},S)\geq\frac{1}{2}c_{H}(\{v\},S)$ total flow from vertices in $S$. Since $A\setminus S_{t}\subseteq V_{H}\setminus(S\cup\{t\})$, the same holds for all $v\in A\setminus S_{t}$. By construction of $H$, we have $c_{H}(\{v\},S)\geq c_{H}(v,s)=\phi\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}(v)+\epsilon\phi\mathbf{d}(v)$ for all $v\in A\setminus S_{t}$. It follows that each vertex $v\in A\setminus S_{t}$ receives at least $\frac{\phi}{2}\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ total flow from vertices in $S$.
We now investigate the effect of restricting the flow $f$ to edges in $G[A\setminus S_{t}]$, starting with removing all edges incident to $S$. Continuing the argument above, removing these edges causes each vertex $v\in A\setminus S_{t}$ to be the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow.
If $x\notin S$, then we now remove the edges incident to $x$. By construction of $H$, each vertex $v\in A$ has an edge to $x$ of capacity $\gamma\mathbf{d}(v)=\frac{\epsilon\phi}{2}\mathbf{d}(v)$. Since the flow $f$ is feasible, there is at most $\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow along the edge $(v,x)$. Removing this edge changes the net flow out of $v$ by at most $\frac{\epsilon\phi}{2}\mathbf{d}(v)$. Since each vertex $v\in A\setminus S_{t}$ is the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}(v)+\frac{\epsilon\phi}{2}\mathbf{d}(v)$ flow before this step, it is the source of at least $\frac{\phi}{2}\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}(v)\geq 0$ flow after this step. In particular, it cannot become a sink. Also, if $v\in A_{t}^{\ell}\setminus S_{t}\subseteq A_{t-1}\setminus A_{t}^{r}$, then it is the source of at least $\frac{\phi}{2}\mathbf{d}(v)$ flow after restriction.
@@
Finally, scaling this restricted flow by $1/(12\phi)$, we conclude that the scaled flow $f$ has congestion $1/(12\phi)$, each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of at least $\mathbf{d}(v)/24$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. ∎
@@ A.3 Analyzing the Potential Decrease
$\deg_{M_{t}}(u)=\mathbf{d}(u)/24$ for all $u\in A_{t}^{\ell}\setminus S_{t}$, and $\deg_{M_{t}}(v)\leq\mathbf{d}(v)$ for all $u\in A_{t}^{r}\setminus S_{t}$.
@@
For each vertex $u\in A_{t}^{\ell}\setminus S_{t}$, its degree in $M_{t}$ equals the total capacity of paths starting at $u$ in the decomposition of $f_{t}$, which is exactly $\mathbf{d}(u)/24$ by Claim A.8. A symmetric argument establishes $\deg_{M_{t}}(v)\leq\mathbf{d}(v)$ for all $u\in A_{t}^{r}\setminus S_{t}$. ∎
@@
We now analyze the potential decrease in the following technical lemma.
@@
$\psi(t)-\psi(t-1)\geq\displaystyle\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}+\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}.$
@@
Subclaim A.18.
$\mathbf{0}\preceq D^{-1/2}LD^{-1/2}\preceq 2I$.
@@
Consider the normalized Laplacian $D_{L}^{-1/2}LD^{-1/2}_{L}$ where $D_{L}$ is the diagonal matrix with value $\deg_{M_{t}}(u)$ on each entry $(u,u)$. It is well-known that the normalized Laplacian has eigenvalues in the range $[0,2]$, so $\mathbf{0}\preceq D_{L}^{-1/2}LD_{L}^{-1/2}\preceq 2I$. Multiplying by $D_{L}^{1/2}$ on both sides gives $\mathbf{0}\preceq L\preceq 2D_{L}$. Claim A.14 implies that $D_{L}\preceq D$, so we obtain $\mathbf{0}\preceq L\preceq 2D_{L}\preceq 2D$. Finally, multiplying by $D^{-1/2}$ on both sides gives the desired $\mathbf{0}\preceq D^{-1/2}LD^{-1/2}\preceq 2I$. ∎
We also need the following fact from linear algebra, whose routine proof we defer to the end of this subsection.
Fact A.19.
For any symmetric matrices $A,B\in\mathbb{R}^{n\times n}$ satisfying $A\succeq\mathbf{0}$ and $-I\preceq B\preceq I$, we have $\textup{Tr}(B^{\top}AB)\leq\textup{Tr}(A)$.
@@
Let $J\in\mathbb{R}^{A\times A}$ denote the all-ones matrix, and for two matrices $A,B\in\mathbb{R}^{A\times A}$, define the Hadamard product $A\bullet B=\sum_{u,v\in A}A(u,v)\cdot B(u,v)=\textup{Tr}(A^{\top}B)$. We first write $\psi(t-1)$ as
@@
$=\sum_{u\in A}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}-\sum_{u\in A\setminus A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$=\sum_{u\in A}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)^{1/2}}-\mathbf{d}(u)^{1/2}\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}-\sum_{u\in A\setminus A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$=(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})-\sum_{u\in A\setminus A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}.$
To bound $\psi(t)$, we use the fact that $\boldsymbol{\mu}_{t}$ minimizes the expression in (2) to get
@@
$\leq\sum_{u\in A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$=\sum_{u\in A}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}-\sum_{u\in A\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$=\sum_{u\in A}\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)^{1/2}}-\mathbf{d}(u)^{1/2}\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}-\sum_{u\in A\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$=(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})-\sum_{u\in A\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}.$
Taking the difference and using the fact that $\mathbf{F}_{t-1}(u)=\mathbf{F}_{t}(u)$ for $u\in A\setminus A_{t}$, and then expanding,
$\psi(t-1)-\psi(t)-\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$
$\geq(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})-(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})$
$=(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1})-(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})\bullet(\widetilde{\mathbf{F}}_{t}-D^{1/2}J\boldsymbol{\mu}_{t-1})$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-2\widetilde{\mathbf{F}}_{t-1}\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}+D^{1/2}J\boldsymbol{\mu}_{t-1}\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}$
$\qquad\quad\,-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}+2\widetilde{\mathbf{F}}_{t}\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}-D^{1/2}J\boldsymbol{\mu}_{t-1}\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}+2(\widetilde{\mathbf{F}}_{t}-\widetilde{\mathbf{F}}_{t-1})\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}$
$=\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}-2\bigg{(}\frac{1}{2}D^{-1/2}LD^{-1/2}\bigg{)}\bullet D^{1/2}J\boldsymbol{\mu}_{t-1}.$
The third term equals $\textup{Tr}\big{(}D^{-1/2}LD^{-1/2}D^{1/2}J\boldsymbol{\mu}_{t-1}\big{)}$, which equals 00 since $LD^{-1/2}D^{1/2}J=LJ$ is the zero matrix. Expanding the first and second terms, we obtain
$\psi(t-1)-\psi(t)-\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}$
$\geq\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}-\widetilde{\mathbf{F}}_{t}\bullet\widetilde{\mathbf{F}}_{t}$
@@
To bound the first term in (5), recall from Subclaim A.18 that $\mathbf{0}\preceq D^{-1/2}LD^{-1/2}\preceq 2I$, which means that $-I\preceq N\preceq I$. Since $\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top}\succeq\mathbf{0}$ and $-I\preceq N\preceq I$, we have $\textup{Tr}(N\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top}N)\leq\textup{Tr}(\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top})$ by Fact A.19, so
@@
$\psi(t-1)-\psi(t)-\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}\geq\frac{1}{2}D^{-1/2}LD^{-1/2}\widetilde{\mathbf{F}}_{t-1}\bullet\widetilde{\mathbf{F}}_{t-1}.$
@@
$=\sum_{u,v\in A}c_{M_{t}}(u,v)\cdot(\mathbbm{1}_{u}-\mathbbm{1}_{v})^{\top}D^{-1/2}\widetilde{\mathbf{F}}_{t-1}\widetilde{\mathbf{F}}_{t-1}^{\top}D^{-1/2}(\mathbbm{1}_{u}-\mathbbm{1}_{v})$
@@
$\psi(t-1)-\psi(t)-\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}\geq\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\cdot\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}.\qed$
@@
$\psi(t-1)-\psi(t)\geq\displaystyle\frac{1}{2}\sum_{u,v\in A}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}+\sum_{u\in A_{t-1}\setminus A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}.$ (6)
@@
$(\mathbf{p}(u)-\mathbf{p}(v))^{2}\leq\frac{C\log n}{n}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2},\quad(\mathbf{p}(u)-\bar{\mu}_{t-1})^{2}\leq\frac{C\log n}{n}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}$ (8)
for all vertices $u,v\in A$ with high probability. By Claim A.14, we have $\deg_{M_{t}}(u)=\mathbf{d}(u)/24$ for all $u\in A_{t}^{\ell}\setminus S_{t}$, and together with statements a to d of Lemma A.5, we have
@@
$=\sum_{u\in A_{t}^{\ell}\setminus S_{t}}\deg_{M_{t}}(u)\cdot(\mathbf{p}(u)-\eta)^{2}$
$=\frac{1}{24}\sum_{u\in A_{t}^{\ell}\setminus S_{t}}\mathbf{d}(u)\cdot(\mathbf{p}(u)-\eta)^{2}$
@@
$=\mathbb{E}\bigg{[}\displaystyle\sum_{\begin{subarray}{c}u\in A_{t}^{\ell}\setminus S_{t}\\ v\in A_{t}^{r}\setminus S_{t}\end{subarray}}c_{M_{t}}(u,v)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}+\sum_{u\in S_{t}\cap A_{t-1}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2}\bigg{]}$
@@
$\geq\mathbb{E}\bigg{[}\frac{n}{216C\log n}\sum_{u\in A_{t}^{\ell}}\mathbf{d}(u)\cdot(\mathbf{p}(u)-\bar{\mu}_{t-1})^{2}\bigg{]}$
@@
$=\frac{n}{80\cdot 216C\log n}\sum_{u\in A_{t-1}}\mathbf{d}(u)\cdot\mathbb{E}[(\mathbf{p}(u)-\bar{\mu}_{t-1})^{2}]$
[…]
$=\frac{1}{80\cdot 216C\log n}\cdot\psi(t-1),$